# FLOPS: Forward Learning with OPtimal Sampling

### v1

@@ 5.1 Finetuning Vision Transformer
model Method ImageNet Caltech101 Food101 Flowers102 CIFAR10/100 EuroSAT ViT-base BP (upper bound) 80.5 92.2 92.4 93.5 96.3 / 90.7 91.9 Mezo (2 queries) 9.3 42.6 37.3 38.2 42.1 / 37.5 39.8 DeepZero ($n$-SPSA) 55.6 78.4 77.8 78.4 85.6 / 70.2 78.8 OPS-SPSA 60.8 87.3 87.6 85.6 89.5 / 82.2 83.2 LR 57.3 85.5 87.6 86.5 87.7 / 80.8 82.9 OPS-LR (Ours) 65.8 88.9 88.4 91.6 93.2 / 88.7 89.3 ViT-Large BP (upper bound) 81.7 93.4 93.3 94.2 98.2 / 93.5 93.4 Mezo (2 queries) 10.7 45.6 38.8 40.5 45.9 / 38.3 41.7 DeepZero ($n$-SPSA) 56.4 79.5 78.9 79.6 86.3 / 73.7 80.1 OPS-SPSA 62.2 88.6 88.1 87.3 91.3 / 83.9 87.6 LR 58.2 87.2 87.7 86.8 88.4 / 81.4 85.7 OPS-LR (Ours) 67.6 90.2 89.5 92.3 95.5 / 90.8 90.1
Using different gradient estimators, we conduct full-parameters fine-tuning for the Vision Transformer (both base and large model). We include DeepZero (SPSA with $n$ queries, denoted as $n$-SPSA), Mezo (SPSA with only $2$ queries), and LR with equally allocated query budgets as baselines in our experiments. For the SPSA-based methods, we include a version with the optimal allocator (OPS-SPSA) to compare with their corresponding baselines. We apply the same settings to the LR, comparing the optimally allocated version (OPS-LR) with the vanilla version. We train the network for 100 epochs with the learning rate of $1\times 10^{-4}$ and the Adam optimizer. The batch size is 64. All the methods in our experiments use the same query budgets, except for Mezo, which uses only 2 queries per data point in accordance with its original memory-efficient settings. As shown in Table 1, OPS-LR outperforms all other methods. In contrast, Mezo with only 1 query performs poorly in the experiment. On the contrary, DeepZero achieves acceptable accuracy compared with Mezo. The phenomenon suggests that an adequate number of queries is crucial for the success of perturbation-based methods since the gradient’s variance must be properly controlled. The OPS-SPSA achieves better accuracy than the DeepZero with an equally allocated budget. Our proposed optimal allocator provides universal acceleration across all perturbation-based methods. Moreover, LR achieves higher accuracy than DeepZero, suggesting that LR is a better gradient estimator than the SPSA in this experiment context. Please refer to the Appendix C.1 for the details and further analysis.
@@ 5.2 Black-box tuning for Model as a Service
Figure 2: illustration paradigm of fine-tuning the prompt for black-box vision language model
We can tune the prompts for vision-language models (VLMs) in a Model-as-a-Service (MaaS) environment using the OPS-LR framework, as shown in Figure 2. We address the challenge of fine-tuning VLMs without access to model internals or gradient information through a scalable black-box optimization approach. We use CLIP (Radford et al., 2021), with embedding dimensions of $D_{T}=512$ for the text encoder and $D_{I}=756$ for the image encoder, as the backbone foundation model. Both the text encoder and the image encoder are kept frozen as black-box modules. Tunable prefix or suffix prompt tokens are injected into the text and image token sequences, respectively: $S_{T}=[cls,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{T}},e_{T}]\in\mathbb{R}^{(1+n_{T}+m_{T})\times D_{T}}$ is the input sequence for the text encoder, $S_{I}=[cls,e_{I},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{I}}]\in\mathbb{R}^{(1+m_{I}+n_{I})\times D_{I}}$ for the image counterpart. Previous researches have suggested that large foundation models have low intrinsic dimensions. Therefore, it is plausible to project the original embedding space onto a low-dimension subspace: $p_{T}=z_{T}\cdot W_{T};W_{T}\in\mathbb{R}^{d_{T}\times D_{T}},z_{T}\in\mathbb{R}^{m_{T}\times d_{T}}$ and $p_{I}=z_{I}\cdot W_{I};W_{I}\in\mathbb{R}^{d_{I}\times D_{I}},z_{I}\in\mathbb{R}^{m_{I}\times d_{I}}$. $W_{T}$ and $W_{I}$ are the projection matrix. The projection matrix is kept frozen after being initialized with a Gaussian distribution. With the projection, our black-box tuning problem is reduced to dimensions of $d_{T}+d_{I}\ll D_{T}+D_{I}$. The optimization problem, using the cross-entropy loss, takes the form:
@@
where $Z$ is the unified formulation of $[z_{T},z_{I}]$. By applying the OPS-LR framework, we can optimize the black-box prompt tuning problems without access to the model structure. We can achieve an accuracy of $69.7\%$ on ImageNet by tuning only a few thousand parameters, surpassing the performance of other black-box methods and Manual Prompting (Radford et al., 2021). Please refer to the Appendix C.2 for more details.
@@ 5.3 Alignment between black-box foundation model
Figure 3: Scalable perturbation-based method can reduce the training time for feature alignment.
Foundation models like GPT (Brown, 2020), LLaMa Touvron et al. (2023), Vicuna (Chiang et al., 2023), and CLIP (Radford et al., 2021) contain rich domain knowledge from the pretraining. Multimodal tasks, such as video captioning (Krishna et al., 2017) and video grounding(Gao et al., 2017), require models to have sufficient cross-domain understanding, and proper alignment between different modalities is essential. Aligning foundation models from a single modality is a more economical approach than pretraining from scratch using multimodal data. It is a common practice to add an adapter between the visual encoder and the Large Language model. The adapter is a linear projector, $g(\cdot)$, that projects the visual embedding to the textual embedding space. Since foundation models contain billions of parameters, the backward pass through the model requires extra memory for storing the computation graph, and recursive computations through such a large pre-trained model are time-consuming. Our OPS-LR framework efficiently bypasses the need for recursive computation to estimate the gradient with a large foundation model in between, as shown in Figure 3(a).
The training paradigm for the multimodal alignment follows an autoregressive style using image-text pairs $<I,T>$. We use a frozen CLIP ViT-L/14 as the visual encoder, denoted as $ViT$. The CLS tokens of the output sequence of $ViT$ are utilized as the global feature for the frame. We denote the visual token after the projection as $Z_{v}=g(ViT(I)^{CLS})$. The visual token is then concatenated with the text tokens sequence: $[Z_{v},t_{1},t_{2},...,t_{m}]$, where $[t_{1},t_{2},...,t_{m}]$ is the text tokens sequence with a length of $m$. Consequently, we can train the adapter on the concatenated sequence with the autoregressive objective of the LLM. We report the wall time for the feature alignment procedure on the LCS-558K dataset (Liu et al., 2024). Our OPS-LR framework can reduce almost half of the training time since the recursive backward computation is skipped. From Figure 3(b), the wall time with or without the allocator is almost the same, which further illustrates that our allocator is lightweight and has minimal impact on runtime. Please refer to the Appendix C.3 for more details.
5.4 Abation study
@@
Impact of different allocators, Gaussian or Bernoulli. We compare different allocators in Figure 4(a), 4(b),4(d), and 4(e). GA controls the allocation with the mean vectors and the covariance matrix. The covariance matrix captures the similarity between different data points. Structurally similar data points possess similar levels of importance. Therefore, the allocation policy for those data points should be correlated. If the covariance matrix is diagonal, the similarity between data points is ignored, which could undermine the performance of the allocation strategy. Compared with the GA, the BA is a simpler strategy. Intuitively, it prunes the queries on unimportant data points to eliminate unnecessary computation. Our ablation study suggests that GA is superior to BA.
Estimation on different components of Transformer. We present the cosine similarity results on the multi-head attention (MHA) and the feed-forward network (FFN) in Figure 4(a), 4(b), 4(c), and 4(f). Estimation on the FFN is easier than on the MHA. In the Transformer block, the MHA precedes the FFN, which may explain this phenomenon. The low cosine similarity on MHA attention of deep layers may lead to the performance gap on large-scale dataset like ImageNet. It is also worth noting that a large number of queries when training Transformer can improve the estimation quality. However, the memory cost of increasing queries is extremely high since the complexity of the attention module is $\mathcal{O}(n^{2})$. Integrating efficient MHA techniques, such as Flash Attention (Dao et al., 2022), is left for future work.
@@ 6 Conclusions
In this paper, we unify different forward learning under the perturbation-based framework. We propose an optimal allocation method to accelerate forward learning by controlling the gradient variance in the mini-batch. Theoretical results show that the allocation can be conducted in a lightweight style and the optimality of the allocation is proved. Compared with other forward learning methods, we achieve the SOTA performance in fine-tuning ViT. With the improved scalability, we apply our method to two black-box scenarios for the foundation model. However, there is still a non-neglectable gap compared with BP. Other techniques of accelerating forward learning are left to future work. Moreover, forward learning can optimize black-box objectives, such as human preference, where taking derivative is infeasible. It is also an interesting future direction.
@@ A.2 Proof of Theorem 1
$LB_{t}$ $=\mathbb{E}_{\bm{A}}\sum_{j=1}^{B}\bigg{[}\big{(}\eta_{t}\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t}))^{\top}\nabla_{\theta}\mathcal{L}_{j}(\theta_{t})-\frac{1}{2}\eta_{t}^{2}L\|\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})\|^{2}\bigg{]}$ (14)
$=\mathbb{E}_{\bm{A}}\sum_{j=1}^{B}\bigg{[}\big{(}\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})\big{)}^{\top}\nabla_{\theta}\mathcal{L}_{j}(\theta_{t})-\frac{1}{2}\eta_{t}^{2}L\|\frac{1}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})\|^{2}\bigg{]},$
@@
$\forall j=1,\cdots,B$, the expected inner product on the RHS of (15)
@@
(17$)=\frac{1}{B^{2}}\bigg{(}\bigg{\|}\sum_{j=1}^{B}\mu_{G}^{j,t}\bigg{\|}^{2}+\sum_{j=1}^{B}\operatorname{Tr}\bigg{(}\frac{\Sigma_{G}^{j,t}}{A_{j}}\bigg{)}\bigg{)}.$ (21)
@@ A.3 Proof of Theorem 2
$\mathbb{E}\big{(}LB^{\ast}_{1:T}-LB_{1:T}^{\mathrm{equal}}\big{)}$ $\geq\mathbb{E}\big{(}LB^{\mathrm{det}}_{1:T}-LB_{1:T}^{\mathrm{equal}}\big{)}=\sum_{t=1}^{\top}\frac{\eta_{t}^{2}L}{2B}(J_{t}(\bm{A}^{\mathrm{equal}})-J_{t}(\bm{A}^{\mathrm{det}}))$ (24)
$\geq\sum_{t=1}^{\top}\frac{\eta_{t}^{2}L}{2BA_{0}}\sum_{j<k}\left(\sqrt{\mathrm{Tr}(\Sigma^{j,t}_{G})}-\sqrt{\mathrm{Tr}(\Sigma^{k,t}_{G})}\right)^{2},$
@@ C.1 Experiments for ViT
ViT-base: The base model has 12 transformer layers. The hidden dimension is 786, with 12 heads for the multi-head attention. The intermediate dimension for feed-forward mlp is 3072. 12 layers are equally divided into three groups from bottom to top. The standard deviation of the injected Gaussian noise for each group is initialized as $1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}$. The query budget for each group is $20$, $60$, and $120$ queries per data. The query budget for the classification head is $60$ queries per data.
ViT-large: The large model has 24 transformer layers. The hidden dimension is 1024, with 16 heads for the multi-head attention. The intermediate dimension for feed-forward mlp is 4096. 24 layers are equally divided into four groups from bottom to top. The standard deviation of the injected Gaussian noise for each group is initialized as $5\times 10^{-6},1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}$. The query budget for each group is $20$, $20$, $60$, and $120$ queries per data. The query budget for the classification head is $60$ queries per data.

### v3

@@ 5.1 Finetuning Vision Transformer
Model Method ImageNet Caltech101 Food101 Flowers102 CIFAR10/100 EuroSAT ViT-base BP (upper bound) 80.5 92.2 92.4 93.5 96.3 / 90.7 91.9 Mezo (2 queries) 9.3 42.6 37.3 38.2 42.1 / 37.5 39.8 Mezo ($n$-SPSA) 54.8 77.2 76.7 80.2 84.4 / 71.5 78.4 DeepZero ($n$-SPSA) 55.6 78.4 77.8 78.4 85.6 / 70.2 78.8 LR 57.3 85.5 87.6 86.5 87.7 / 80.8 82.9 OPS-SPSA (Ours) 60.8 87.3 87.6 85.6 89.5 / 82.2 83.2 OPS-LR (Ours) 65.8 88.9 88.4 91.6 93.2 / 88.7 89.3 ViT-Large BP (upper bound) 81.7 93.4 93.3 94.2 98.2 / 93.5 93.4 Mezo (2 queries) 10.7 45.6 38.8 40.5 45.9 / 38.3 41.7 Mezo ($n$-SPSA) 56.7 78.8 78.4 78.5 86.7 / 74.3 79.2 DeepZero ($n$-SPSA) 56.4 79.5 78.9 79.6 86.3 / 73.7 80.1 LR 58.2 87.2 87.7 86.8 88.4 / 81.4 85.7 OPS-SPSA (Ours) 62.2 88.6 88.1 87.3 91.3 / 83.9 87.6 OPS-LR (Ours) 67.6 90.2 89.5 92.3 95.5 / 90.8 90.1
Using different gradient estimators, we conduct full-parameters fine-tuning for the Vision Transformer (both base and large model). We include DeepZero (SPSA with $n$ queries, denoted as $n$-SPSA), Mezo (SPSA with only $2$ queries), and LR with equally allocated query budgets as baselines in our experiments. For the SPSA-based methods, we include a version with the optimal allocator (OPS-SPSA) to compare with their corresponding baselines. We apply the same settings to the LR, comparing the optimally allocated version (OPS-LR) with the vanilla version.
We train the network for 10 epochs with the learning rate of $1\times 10^{-4}$ and the Adam optimizer. The batch size is 128. All the methods in our experiments use the same query budgets, except for Mezo, which uses only 2 queries per data point in accordance with its original memory-efficient settings. As shown in Table 1, OPS-LR outperforms all other methods. In contrast, Mezo with only 1 query performs poorly in the experiment. On the contrary, DeepZero achieves acceptable accuracy compared with Mezo. The phenomenon suggests that an adequate number of queries is crucial for the success of perturbation-based methods since the gradient’s variance must be properly controlled. The OPS-SPSA achieves better accuracy than the DeepZero with an equally allocated budget. Our proposed optimal allocator provides universal acceleration across all perturbation-based methods. Moreover, LR achieves higher accuracy than DeepZero, suggesting that LR is a better gradient estimator than the SPSA in this experiment context. Please refer to the Appendix C.1 for the details and further analysis.
@@ 5.2 Black-box tuning for Model as a Service
Figure 2: illustration paradigm of fine-tuning the prompt for black-box vision language model.
We can tune the prompts for vision-language models (VLMs) in a Model-as-a-Service (MaaS) environment using the OPS-LR framework, as shown in Figure 2. We address the challenge of fine-tuning VLMs without access to model internals or gradient information through a scalable black-box optimization approach. We use CLIP (Radford et al., 2021), with embedding dimensions of $D_{T}=512$ for the text encoder and $D_{I}=756$ for the image encoder, as the backbone foundation model. Both the text encoder and the image encoder are kept frozen as black-box modules.
Tunable prefix or suffix prompt tokens are injected into the text and image token sequences, respectively: $S_{T}=[cls,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{T}},e_{T}]\in\mathbb{R}^{(1+n_{T}+m_{T})\times D_{T}}$ is the input sequence for the text encoder, $S_{I}=[cls,e_{I},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{I}}]\in\mathbb{R}^{(1+m_{I}+n_{I})\times D_{I}}$ for the image counterpart. Previous researches have suggested that large foundation models have low intrinsic dimensions. Therefore, it is plausible to project the original embedding space onto a low-dimension subspace: $p_{T}=z_{T}\cdot W_{T};W_{T}\in\mathbb{R}^{d_{T}\times D_{T}},z_{T}\in\mathbb{R}^{m_{T}\times d_{T}}$ and $p_{I}=z_{I}\cdot W_{I};W_{I}\in\mathbb{R}^{d_{I}\times D_{I}},z_{I}\in\mathbb{R}^{m_{I}\times d_{I}}$. $W_{T}$ and $W_{I}$ are the projection matrix. The projection matrix is kept frozen after being initialized with a Gaussian distribution. With the projection, our black-box tuning problem is reduced to dimensions of $d_{T}+d_{I}\ll D_{T}+D_{I}$. The optimization problem, using the cross-entropy loss, takes the form:
@@
where $Z$ is the unified formulation of $[z_{T},z_{I}]$. By applying the OPS-LR framework, we can optimize the black-box prompt tuning problems without access to the model structure. We can achieve an accuracy of $69.7\%$ on ImageNet by tuning only a few thousand parameters, surpassing the performance of other black-box methods and Manual Prompting (Radford et al., 2021). The performance is shown in Table 2. Please refer to the Appendix C.2 for more details.
@@ 5.3 Alignment between black-box foundation model
Foundation models like GPT (Brown, 2020), LLaMa Touvron et al. (2023), Vicuna (Chiang et al., 2023), and CLIP (Radford et al., 2021) contain rich domain knowledge from the pretraining. Multimodal tasks, such as video captioning (Krishna et al., 2017) and video grounding(Gao et al., 2017), require models to have sufficient cross-domain understanding, and proper alignment between different modalities is essential.
Figure 3: illustration alignment between foundation model for video understanding.
Aligning foundation models from a single modality is a more economical approach than pretraining from scratch using multimodal data. It is a common practice to add an adapter between the visual encoder and the Large Language model. The adapter is a linear projector, $g(\cdot)$, that projects the visual embedding to the textual embedding space. Since foundation models contain billions of parameters, the backward pass through the model requires extra memory for storing the computation graph, and recursive computations through such a large pre-trained model are time-consuming. Our OPS-LR framework efficiently bypasses the need for recursive computation to estimate the gradient with a large foundation model in between, as shown in Figure 3.
The training paradigm for the multimodal alignment follows an autoregressive style using image-text pairs $<I,T>$. We use a frozen CLIP ViT-L/14 as the visual encoder, denoted as $ViT$. The CLS tokens of the output sequence of $ViT$ are utilized as the global feature for the frame. We denote the visual token after the projection as $Z_{v}=g(ViT(I)^{CLS})$. The visual token is then concatenated with the text tokens sequence: $[Z_{v},t_{1},t_{2},...,t_{m}]$, where $[t_{1},t_{2},...,t_{m}]$ is the text tokens sequence with a length of $m$. Consequently, we can train the adapter on the concatenated sequence with the autoregressive objective of the LLM. We report the wall time for the feature alignment procedure on the LCS-558K dataset (Liu et al., 2024). Our OPS-LR framework can reduce almost half of the training time since the recursive backward computation is skipped. From Figure 5, the wall time with or without the allocator is almost the same, which further illustrates that our allocator is lightweight and has minimal impact on runtime. Please refer to the Appendix C.3 for more details.
5.4 Ablation study
@@
Figure 5: Wall clock time for feature alignment between different methods.
@@
Impact of different allocators, Gaussian or Bernoulli. We compare different allocators in Figures 4(4(a)), (4(b)), (4(d)), and (4(e)). GA controls the allocation with the mean vectors and the covariance matrix. The covariance matrix captures the similarity between different data points. Structurally similar data points possess similar levels of importance. Therefore, the allocation policy for those data points should be correlated. If the covariance matrix is diagonal, the similarity between data points is ignored, which could undermine the performance of the allocation strategy. Compared with the GA, the BA is a simpler strategy. Intuitively, it prunes the queries on unimportant data points to eliminate unnecessary computation. Our ablation study suggests that GA is superior to BA.
Estimation on different components of Transformer. We present the cosine similarity results on the multi-head attention (MHA) and the feed-forward network (FFN) in Figures 4(4(a))-(4(c)) and (4(f)). Estimation on the FFN is easier than on the MHA. In the Transformer block, the MHA precedes the FFN, which may explain this phenomenon. The low cosine similarity on MHA attention of deep layers may lead to the performance gap on large-scale dataset like ImageNet. It is also worth noting that a large number of queries when training Transformer can improve the estimation quality. However, the memory cost of increasing queries is extremely high since the complexity of the attention module is $\mathcal{O}(n^{2})$. Integrating efficient MHA techniques, such as Flash Attention (Dao et al., 2022), is left for future work.
@@ 6 Conclusions
In this paper, we present a unified perspective on forward learning through a perturbation-based framework and propose an optimal query allocation strategy to enhance its efficiency. By leveraging a likelihood ratio-based objective and a reparameterized Gaussian allocator, our method adaptively distributes queries based on data-specific gradient estimation difficulty, ensuring an optimal balance between computational cost and estimation accuracy. Theoretical analysis guarantees the optimality of our allocation strategy, while extensive experiments demonstrate its effectiveness in fine-tuning Vision Transformers and addressing real-world black-box optimization problems, such as prompt tuning and multimodal model alignment. Our results show that the proposed method significantly improves the scalability of forward learning, achieving state-of-the-art performance compared to existing approaches. While our method significantly improves scalability, closing the gap with backpropagation remains an open challenge for future research.
@@ A.2 Proof of Theorem 1
$LB_{t}$ $=\mathbb{E}\sum_{j=1}^{B}\bigg{[}\big{(}\eta_{t}\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t}))^{\top}\nabla_{\theta}\mathcal{L}_{j}(\theta_{t})-\frac{1}{2}\eta_{t}^{2}L\|\nabla_{\theta}\hat{\mathcal{L}}(\theta_{t})\|^{2}\bigg{]}$ (16)
$=\mathbb{E}\sum_{j=1}^{B}\bigg{[}\big{(}\frac{\eta_{t}}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})\big{)}^{\top}\nabla_{\theta}\mathcal{L}_{j}(\theta_{t})-\frac{1}{2}\eta_{t}^{2}L\|\frac{1}{B}\sum_{j=1}^{B}\nabla_{\theta}\hat{\mathcal{L}}_{j}(\theta_{t})\|^{2}\bigg{]},$
@@
$\forall j=1,\cdots,B$, the expected inner product on the LHS of (17)
@@
(19$)=\frac{1}{B^{2}}\bigg{(}\bigg{\|}\sum_{j=1}^{B}\mu_{G}^{j,t}\bigg{\|}^{2}+\sum_{j=1}^{B}\operatorname{Tr}\bigg{(}\frac{\Sigma_{G}^{j,t}}{A_{j}}\bigg{)}\bigg{)}.$ (23)
@@ A.3 Proof of Theorem 2
$\mathbb{E}\big{(}LB^{\ast}_{1:T}-LB_{1:T}^{\mathrm{equal}}\big{)}$ $\geq\mathbb{E}\big{(}LB^{\mathrm{det}}_{1:T}-LB_{1:T}^{\mathrm{equal}}\big{)}=\sum_{t=1}^{T}\frac{\eta_{t}^{2}L}{2B}(J_{t}(\bm{A}^{\mathrm{equal}})-J_{t}(\bm{A}^{\mathrm{det}}))$ (26)
$\geq\sum_{t=1}^{T}\frac{\eta_{t}^{2}L}{2BA_{0}}\sum_{j<k}\left(\sqrt{\mathrm{Tr}(\Sigma^{j,t}_{G})}-\sqrt{\mathrm{Tr}(\Sigma^{k,t}_{G})}\right)^{2},$
@@ C.1 Experiments for ViT
ViT-base: The base model has 12 transformer layers. The hidden dimension is 786, with 12 heads for the multi-head attention. The intermediate dimension for feed-forward mlp is 3072. 12 layers are equally divided into three groups from bottom to top. The standard deviation of the injected Gaussian noise for each group is initialized as $1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}$. The query budget for each group is $20$ queries per data. The query budget for the classification head is also $20$ queries per data.
ViT-large: The large model has 24 transformer layers. The hidden dimension is 1024, with 16 heads for the multi-head attention. The intermediate dimension for feed-forward mlp is 4096. 24 layers are equally divided into four groups from bottom to top. The standard deviation of the injected Gaussian noise for each group is initialized as $5\times 10^{-6},1\times 10^{-5},1\times 10^{-4},1\times 10^{-3}$. The query budget for each group is $20$ queries per data. The query budget for the classification head is $20$ queries per data.